# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v4

@@ Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared representation space via contrastive learning on large-scale image-text pairs. Its effectiveness primarily stems from the use of natural language as rich supervision. Motivated by the remarkable advancements in large language models (LLMs), this work explores how LLMs’ superior text understanding and extensive open-world knowledge can enhance CLIP’s capability, especially for processing longer and more complex image captions. We propose an efficient post-training strategy that integrates LLMs into pretrained CLIP. To address the challenge posed by the autoregressive nature of LLMs, we introduce a caption-to-caption contrastive fine-tuning framework, significantly enhancing the discriminative quality of LLM outputs. Extensive experiments demonstrate that our approach outperforms LoRA-based methods, achieving nearly fourfold faster training with superior performance. Furthermore, we validate substantial improvements over state-of-the-art models such as CLIP, EVA02, and SigLip2 across various zero-shot multimodal retrieval tasks, cross-lingual retrieval tasks, and multimodal language model pretraining.
@@ 1 Introduction
Figure 1: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
CLIP [28] is one of the most influential multimodal foundational models today. What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [6] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
@@
This paper successfully explores a efficient method of leveraging LLM capabilities in CLIP pretraining. LLMs, as text generation models, are primarily pretrained to capture the distribution of natural language, thus their outputs naturally lie close to textual spaces. However, what CLIP pretraining actually requires is a robust text embedding model. Previous works such as LLM2Vec [2] and NV-Embed-v2 [17] have demonstrated that fine-tuning LLMs significantly enhances their ability as text encoders. Nevertheless, it remains unclear how directly applying LLMs affects CLIP performance. Our preliminary results indicate that directly integrating LLMs into CLIP pretraining can be detrimental because LLMs struggle to differentiate effectively between image descriptions, as illustrated by example cases in Figure 3. Moreover, as shown in Table 1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs) and observed that large language models like Llama 3 8B achieve very low accuracy, barely distinguishing between different captions of the same image.
Language Model Top 1 Acc.
CLIP-L/14 25.2
EVA02-L/14 27.1
Llama3-8B 5.2
Llama3.2-1B 5.6
Llama3-8B-CC 29.5
Llama3.2-1B-CC 29.4
Table 1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO [22] 5K test set.
@@
Figure 3: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
We designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of LLM using image captioning dataset as in  Figure 2. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [8, 2] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [3] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 1, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
The main contributions of this paper are:
  1. 1.
    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
  2. 2.
    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space. This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
  3. 3.
    We successfully verified that LLM2CLIP can substantially enhance the performance of CLIP and EVA02 across multiple tasks, including short-text retrieval, long-text retrieval, cross-language retrieval, and multimodal large language model pretraining, through a simple post-training strategy. Remarkably, LLM2CLIP can further improve the short-text retrieval accuracy of state-of-the-art large-scale pretrained models like SigLip2 [33] by +1 and +1.9 points, achieve impressive gains of +14.8 and +15.8 points in long-text retrieval, and deliver significant improvements of +11.9 and +15.2 points in multilingual tasks.
@@ 2 Related Works
CLIP meets Stronger Language Models. Several works have explored the integration of LLMs into CLIP. JinaCLIP [14] employed Jina-embeddings-v2 [9] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [30], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [12] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [18] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [10] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
CLIP meets Longer Captions. It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [34] hired annotators to expand captions based on SAM [13]’s target hints. LaCLIP [7] adopted ChatGPT, Bard, and human rewriting. DreamLIP [41] leveraged ShareCaptioner [3] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [21] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [34], splitting captions into multiple segments [7, 41], or finetuning model’s positional encoding to support longer token inputs [39]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
Methods Resolution Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
@@
ALIGN [11] 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP [19] 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Long-CLIP [39] 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
jina-clip-v2 [15] 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
SigLIP [38] 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
MetaCLIP [36] 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
CLIP [28] 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
@@
Long-CLIP [39] 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
MetaCLIP [36] 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
SAIL-L-NV2 [40] 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
CLIP [28] 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
@@
SigLIP [38] 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
SigLIP2 [33] 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
@@
Table 2: Systematic comparison of model performance on multiple datasets. The light blue row indicate LLM2CLIP post trained version.
Models Flickr-CN [16] COCO-CN [20] XM3600 [32]

### v5

# LLM2CLIP: Powerful Language Model Unlocks Richer Cross-Modality Representation

@@ Abstract
CLIP is a seminal multimodal model that maps images and text into a shared representation space by contrastive learning on billions of image–caption pairs. Inspired by the rapid progress of large language models (LLMs), we investigate how the superior linguistic understanding and broad world knowledge of LLMs can further strengthen CLIP—particularly in handling long, complex captions. We introduce an efficient fine-tuning framework that embeds an LLM into a pretrained CLIP while incurring almost the same training cost as regular CLIP fine-tuning. Our method first “embedding-izes” the LLM for the CLIP setting, then couples it to the pretrained CLIP vision encoder through a lightweight adaptor trained on only a few million image–caption pairs. With this strategy we achieve large performance gains—without large-scale retraining—over state-of-the-art CLIP variants such as EVA02 and SigLIP-2. The LLM-enhanced CLIP delivers consistent improvements across a wide spectrum of downstream tasks, including linear-probe classification, zero-shot image–text retrieval with both short and long captions (in English and other languages), zero-shot/supervised image segmentation, object detection, and used as tokenizer for multimodal large-model benchmarks. Code & Models: https://aka.ms/llm2clip.
@@ 1 Introduction
Figure 1: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer textual supervision.
CLIP (Radford et al. 2021; Tschannen et al. 2025) has emerged as one of the most influential cross-modal foundation models in recent years. Trained on hundreds of millions to tens of billions of image–text pairs via contrastive pre-training, it embeds vision and language into a shared representation space. As a retriever, CLIP underpins zero-shot classification, detection, segmentation, and—most notably—image–text retrieval. As a feature extractor, it supports a broad spectrum of cross-modal applications, from image/video understanding to text-to-image and text-to-video generation. For instance, Multi-modality large language models such as LLaVA (Li et al. 2024a) and Qwen-VL (Bai et al. 2023) rely on CLIP’s visual features, while image/video generative model like Stable-Diffusion-3 (Esser et al. 2024) and Wan (Kong et al. 2024) leverage its text encoder. Nevertheless, as we push toward broader generality and higher task complexity, the representational capacity inherited from CLIP’s original paradigm is gradually becoming inadequate.
[…]
@@
Bringing LLMs into the CLIP framework, however, raises two key challenges: 1. Feature separability. Vanilla LLM embeddings are not sufficiently discriminative for contrastive training; existing “LLM-as-embedding” recipes target pure-text tasks. We therefore devise a CLIP-specific embeddingization strategy for LLMs that yields far more separable caption features. 2. Training cost. CLIP pre-training is already expensive; naively fine-tuning an LLM jointly would be prohibitive. We propose a lightweight fine-tuning procedure that injects LLM power into a pretrained CLIP at almost no extra cost.
This paper introduces LLM2CLIP, an efficient fine–tuning framework that augments the feature space of a pre-trained CLIP with an embedding-tuned LLM, thereby importing LLM capability at very low cost. Injecting a vanilla LLM into CLIP is problematic: as shown by the cases in Figure 2, a COCO test reveals that raw LLM embeddings have poor separability for image captions. To remedy this, we perform caption-contrastive(CC) fine-tuning on the LLM with a set of high-quality caption datasets, revisiting several design choices—e.g. employing average pooling to aggregate token features, enabling bidirectional attention, and training with a supervised SimCSE contrastive loss.
Another key challenge when introducing LLMs into CLIP is the presence of two text encoders—the original CLIP text encoder and the newly added LLM—which were never aligned during pre-training. We conduct experiments to explore efficient integration of the LLM into the CLIP architecture. As shown in Figure 1, we adopt a cost-effective solution: freeze all LLM gradients, treat its sentence embeddings as fixed features, and append a small learnable adaptor trained with the CLIP visual encoder. The original CLIP text encoder is discarded during both training and inference. Ablation studies indicate that alternative strategies—combining CLIP and LLM embeddings via triplet contrastive losses or concatenating features—provide marginal or negative gains compared to simply replacing the CLIP text encoder. This minimal design yields an LLM-enhanced cross-modal space while keeping compute cost nearly identical to standard CLIP fine-tuning.
Empirical results confirm that LLM2CLIP yields substantial improvements—even with only a few million training examples—boosting the performance of the original CLIP across a variety of downstream tasks. Our contributions are as follows:
  • We empirically demonstrate that injecting LLM capability into CLIP brings significant performance gains.
  • We design (i) a caption-contrastive fine-tuning recipe that turns an LLM into a effective embedding model for CLIP, and (ii) a efficient fine-tuning method that couples this embedding with a pretrained CLIP.
  • Extensive experiments reveal that the resulting LLM-enhanced CLIP markedly improves several state-of-the-art models—including EVA02 and SigLIP-2—on a wide spectrum of multimodal benchmarks: short/long-text and cross-lingual image retrieval, zero-shot classification, detection, segmentation, and even as the visual encoder inside LLaVA-1.5. Remarkably, LLM2CLIP lifts SigLIP-2’s short-caption retrieval by +1.0/+1.9, long-caption retrieval by +14.8/+15.8, and multilingual tasks by +11.9/+15.2.
@@
Figure 2: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often unrelated.
@@ 2 Related Works
CLIP meets Stronger Language Models. Several works have explored the integration of LLMs into CLIP. JinaCLIP (Koukounas et al. 2024a) employed Jina-embeddings-v2 (Günther et al. 2023) as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP (Sun et al. 2023a), its text encoder is far behind ours, limiting the potential benefits from LLMs. MATE (Jang et al. 2024) designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. This paper presents the first comprehensive study of an efficient strategy for incorporating LLMs into CLIP fine-tuning, thereby boosting performance.
CLIP meets Longer Captions. CLIP’s text embedding is widely recognized as coarse and limited to 77 tokens. Many works have attempted to extend caption length and retrain CLIP accordingly, including DCI (Urbanek et al. 2024) with human annotation, LaCLIP (Fan et al. 2024) using ChatGPT and Bard, DreamLIP (Zheng et al. 2024) leveraging ShareCaptioner (Chen et al. 2023), and Recap-DataComp-1B (Li et al. 2024b) using Llama3-trained LLAVA1.5. To handle longer captions, these methods employ workarounds such as summarization (Urbanek et al. 2024), segmentation (Fan et al. 2024; Zheng et al. 2024), or positional encoding finetuning (Zhang et al. 2024a). In comparison, LLM2CLIP leverages LLMs as the text encoder, enabling comprehensive understanding of long and dense captions while utilizing LLMs’ open-world knowledge.
Method Res Flickr COCO SG4V Urban DOCCI Avg
@@
ALIGN (Jia et al. 2021) 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP (Li et al. 2022a) 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Long-CLIP (Zhang et al. 2024a) 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
jina-clip-v2 (Koukounas et al. 2024b) 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
SigLIP (Zhai et al. 2023) 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
MetaCLIP (Xu et al. 2023) 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
CLIP (Radford et al. 2021) 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
@@
Long-CLIP (Zhang et al. 2024a) 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
MetaCLIP (Xu et al. 2023) 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
L-NV2 (Zhang et al. 2025) 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
CLIP (Radford et al. 2021) 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
@@
SigLIP (Zhai et al. 2023) 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
SigLIP2 (Tschannen et al. 2025) 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
@@
Table 1: Systematic comparison of model performance on multiple datasets. * means reproduced by ourselves.
3 Methods
3.1 Stage 1: LLM Caption Contrastive Fine-tuning
Traditionally, an LLM’s output layer functions as a classification head that produces discrete text tokens. Recent studies—such as LLM2Vec (BehnamGhader et al. 2024) and NV-EMBED-v2 (Lee et al. 2024)—have shown that, with additional design tweaks, an LLM can be repurposed as a text-embedding model. Yet these efforts do not account for CLIP’s cross-modal pre-training regime. Here, we advance the discussion from three fronts—model architecture, training methods and training data—explicitly targeting the seamless integration of LLMs with CLIP.
Model Architecture:
  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability. We experimented with three adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs and (3) without using adaptor.
Training Methods:
  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens. Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention. While effective in pure text tasks (as shown by LLM2Vec), we find out that MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
Training Data:
To leverage LLM embeddings for CLIP fine-tuning, we fine-tune the LLM on caption corpora extracted from image–text datasets, aligning its embedding space with downstream multimodal tasks. Experiments use the Dreamlip (Zheng et al. 2024) captions, which provide multiple captions per image, suitable for supervised textual contrastive learning. To preserve general language understanding, we also incorporate 1.5M pure-text pairs from Echo Embeddings (Springer et al. 2024) into a mixed contrastive training setup.
3.2 Stage 2: LLM2CLIP Post Fine-tuning
We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining. By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective fine-tuning step. CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model. Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and finding an effective approach to enhance CLIP performance is critical. Here, we also elaborate on model architecture, training methods and training data.
Model Architecture: Training CLIP typically requires large-scale datasets and substantial batch sizes, resulting in considerable computational costs. Therefore, we employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this stage. By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge. Given the large size of the LLM, we compare two strategies: (1) fine-tuning the LLM with LoRA, and (2) freezing LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain sentence token features. Method (2) offers clear advantages: it completely eliminates LLM gradient updates, significantly reducing GPU memory usage. Moreover, since the adaptor is placed after the LLM, text features can be precomputed and stored via offline inference, reducing inference overhead from multiple epochs to a single pass. As a result, this approach avoids loading the LLM into GPU memory during training, accelerating experimentation and reducing training time and memory footprint. By default, we adopt method (2) with an adaptor of four linear blocks. Detailed efficiency analysis is provided in Section 4.4.
Training Methods: After introducing an LLM into CLIP, we are left with two text encoders. How should they be used most effectively? We tested several approaches: (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning. (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders. (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b). (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder. Our extensive experiments led us to adopt method (a) for its simplicity and effectiveness.
Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions. Datasets such as DreamLIP and Recaption, which leverage Share-Captioner (Chen et al. 2023), InstructBLIP (Dai et al. 2023) and LLaVA-1.5 (Liu et al. 2023a) to generate multiple dense captions per image, can be effectively utilized. We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions. Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
4 Experiments
4.1 Default Experimental Settings
Default Settings for Stage-1 We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention. No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding. MNTP is not used by default.
Default Settings for Stage-2 After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA-3.1-8B to perform cross-modal contrastive learning fine-tuning with the vision encoder. By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM. The design of the Adaptor is the same as FuseMix (Vouitsis et al. 2023), using a simple inverted bottleneck MLP architecture. By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M. For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset. For the 3M setting, we only use the CC 3M subset.
4.2 System Comparison
In the system comparison, we used the default settings from  Section 4.1. However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
Zero-Shot English Text-Image Retrieval For short-text retrieval, we used the MSCOCO (Lin et al. 2014) 5K test set and the Flickr (Young et al. 2014) 1K test set. For long-text retrieval, we employed a 1K subset of ShareGPT4V (Chen et al. 2023) and the Urban1K (Zhang et al. 2024a) dataset from LongCLIP (Zhang et al. 2024a), along with the DOCCI (Onoe et al. 2024) dataset.
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU (Ordonez et al. 2011), and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation. To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval. In  Table 1, we compared OpenAI’s CLIP, EVA02, and SigLIP2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoViT/14. Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02. For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8. Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed. Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts. Notably, LLM2CLIP even surpasses InternVL—which contains 6 B trainable parameters—and VLM2Vec, whose pipeline fine-tunes the entire LLM with LoRA.
Models Flickr-CN COCO-CN XM3600